# Liquid-Graph Time-Constant Network for Multi-Agent Systems Control

### v1

@@ I INTRODUCTION
In recent years, the interest in multi-agent systems (MAS) has significantly increased [1]. The effective coordination of multiple agents necessitates an understanding of group dynamics. Particularly in multi-robot scenarios, generating individual robot motions relies not only on local sensing but also on information about the group’s state, often obtained through communication with a limited subset of nearby team members [2]. As a result, communication becomes a crucial element in achieving distributed solutions for MAS.
Neural networks have become prominent in data-driven control applications due to their outstanding approximation capabilities [3]. In distributed control, neural networks offer particular advantages by efficiently approximating complex distributed policies without the need for intricate optimizations and designs. The literature presents numerous data-driven approaches, with a focus on those rooted in reinforcement learning [4, 5]. However, these approaches operate on local sensing data without inter-agent communication. On the other hand, distributed machine learning employs communication to partition the learning process across multiple machines contributing to group knowledge [6].
@@
In this study, we introduce a continuous graph ordinary differential equation (ODE) network called Liquid-Graph Time-Constant (LGTC) network, inspired by its single-agent counterpart, the Liquid Time Constant (LTC) network [24]. The LTC network demonstrates a strong representation learning capability, empowering agents to extrapolate and make inferences even in unfamiliar scenarios. Additionally, its closed-form version [25] eliminates the need to solve the initial value problem of the ODE, a requirement in other ODE networks [26, 27, 28]. In the existing literature, there are few attempts to make graph ODEs for modelling dynamic network systems [29], predicting traffic flow [30], or for sequential recommendations [31]. These models typically employ an auto-regressive network over a vector field modelled by a Graph Neural Network (GNN), initializing the internal state with the previous layer’s output. In contrast, the proposed LGTC aims to introduce a novel model that learns a dynamic (or “liquid”) time constant, influenced by both input data and the internal state communicated over the agent graph. This dynamic time constant enables each element of the state to capture specific dynamics, thereby enhancing the model’s predictive capabilities.
@@ II PRELIMINARIES
where $\bm{s}_{i}$ is the $i$-th row of $\bm{S}$. The signal manipulation can be operated by means of any graph shift operator $\bm{S}\subseteq\mathcal{S}$, e.g., Laplacian, adjacency matrix, weighted Laplacian , which respects the sparsity pattern of the graph. Later, in Sect. V, we will use the Laplacian as support matrix, as it is commonly used in distributed control. However, the proposed techniques do not assume the use of a specific support matrix.
Performing $k$ repeated applications of $\bm{S}$ on the same signal represents the aggregation of the $k$-hop neighbourhood information. In analogy with traditional signal processing, this property can be used to define a linear graph filtering [32] that processes the multi features signal $\bm{x}\in\mathbb{R}^{N\times G}$ with $G$ features:
$\small H_{\bm{S}}(\bm{x})=\sum_{k=0}^{K}\bm{S}^{k}\bm{x}\bm{H_{k}}$ (2)
where the weights $\bm{H_{k}}\in\mathbb{R}^{G\times F}$ define the output of the filter. Note that $\bm{S}^{k}=\bm{S}(\bm{S}^{k-1})$, so that it can be computed locally with repeated 1-hop communications between a node and its neighbors. Hence, the computation of $H_{\bm{S}}$ is distributed on each node.
@@ II-B Gated Graph Neural Network
Forgetting factors can be applied to mitigate this problem, reducing the influence of past or new signal on the state. A Gated Graph Neural Network (GGNN) [23] is a recurrent Graph Neural Network that uses a gating mechanism to control how the past information influences the update of the GNN states. We can add a state and an input gate, $\bm{\hat{q}},\bm{\tilde{q}}\in Q\subseteq[0,1]^{N\times F}$, that are multiplied via the Hadamard product $\circ$ by the state and the inputs of the network, respectively. These two gates regulate how much the past information and the input are used to update the network’s internal state. GGNNs admit the following state-space representation [36],
@@
We identify the induced $\infty$-norm as $||\cdot||_{\infty}$. We used the following notation for the filters in (4):
@@
where $K$ is the filters length. Then, in light of assumption (1) and knowing that $||\bm{x}||_{\infty}\leq 1$, each gate feature $q_{i}$ satisfies:
$|\hat{q}_{i}|$ $\leq\sigma(||S_{0,K}||_{\infty}(||\hat{A}_{0,K}||_{\infty}+||\hat{B}_{0,K}||_{\infty})+||\hat{\bm{b}}||_{\infty})\triangleq\sigma_{\hat{q}}.$ (6)
@@ III LIQUID-GRAPH TIME CONSTANT NETWORK
In the following results on the system, we used the vector operator $\bm{X}|_{:}$ that rearranges the elements of matrix $\bm{X}$ in a vector. For the system in (8), the following lemma holds
@@
$\begin{split}\dot{V}&=-\bm{x}|_{:}^{T}\operatorname*{diag}(\bm{b}|_{:})\bm{x}|_{:}-\bm{x}|_{:}^{T}\operatorname*{diag}(\bm{f}|_{:})\bm{x}|_{:}\\ &-\bm{x}|_{:}^{T}\sum_{k=1}^{K}\bm{A_{k}}^{T}\otimes\bm{S}^{k}\bm{x}|_{:}+\bm{x}|_{:}^{T}\operatorname*{diag}(\bm{f}|_{:})\sigma_{c}(B_{S}(\bm{u}))|_{:}\end{split}$
@@
$\small\dot{V}<-\bm{x}|_{:}^{T}\operatorname*{diag}(\bm{f}|_{:})\bm{x}|_{:}+\bm{x}|_{:}^{T}\operatorname*{diag}(\bm{f}|_{:})\sigma_{c}(B_{S}(\bm{u}))|_{:}$
When $x_{i}>1$, $\dot{V_{i}}<0$, because the term $\bm{x}|_{:}^{T}\operatorname*{diag}(\bm{f}|_{:})\sigma_{c}(B_{S}(\bm{u}))|_{:}\leq\bm{1}^{T}\operatorname*{diag}(\bm{f}|_{:})\bm{1}$ and reminding that $f_{i}>0$. The proof is completed by noting that $\dot{V_{i}}<0$ when $x_{i}<-1$ with a similar reasoning. ∎
We named the vector field in eq. (8) with $F(x,u,S,t):\mathcal{X}\times\mathcal{U}\times\mathcal{S}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{N\times F}$ and its jacobian with $D_{x}F=\frac{\partial F|_{:}}{\partial\bm{x}|_{:}}$. Giving the induced infinite log-norm definition $\mu_{\infty}(\bm{X})=\max_{i}(x_{ii}+\sum_{j=1,j\neq 1}^{n}|x_{ij}|)$ [37], the followings results are known [15]:
@@
The vector field $F$ is strongly infinitesimally contractivng if:
@@
For the neural network 8, we provide the following theorem
@@
$\small\begin{split}\sup\limits_{x}\mu_{\infty}(D_{x}F)=&\sup\limits_{x}\mu_{\infty}(-\operatorname*{diag}(\bm{b|_{:}})-\operatorname*{diag}(f|_{:})\\ -\sum_{k=1}^{K}\bm{A_{k}}^{T}\otimes\bm{S}^{k}+\operatorname*{diag}(&\sigma_{u}|_{:}-x|_{:})\operatorname*{diag}(Df)\sum_{k=0}^{K}\bm{\hat{A_{k}}}^{T}\otimes\bm{S}^{k})\\ \leq-||\bm{b}||_{\infty}+\mu_{\infty}(&-\sum_{k=1}^{K}\bm{A_{k}}^{T}\otimes\bm{S}^{k})+\sup\limits_{x}-||f|_{:}||_{\infty}\\ +\mu_{\infty}(\operatorname*{diag}(&\sigma_{u}|_{:}-x|_{:})\operatorname*{diag}(Df)\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}\end{split}$
where $D_{x}f\in[0,1]^{NF}$ is the derivative of $f$ in $\bm{x}|_{:}$. The last inequality follows by $\mu_{\infty}(\bm{X})=||\bm{X}||_{\infty}$ when $\bm{X}>0$ and by the fact that $\bm{b}|_{:}\geq 0$ and $f|_{:}\geq 0$. Moreover, we can note that $-||f|_{:}||_{\infty}\leq-||\rho(\hat{A}_{S}(\bm{x})+\bm{b_{x}})|_{:}||_{\infty}$. Giving $\sigma_{u}|_{:}\in[-1,1]^{NF}$ and $-\mu_{\infty}(\bm{A_{k}}^{T}\otimes\bm{S}^{k})\leq 0$, by lemma 1, $\bm{x}|_{:}\in[-1,1]^{NF}$. Therefore, we get
@@
From the statement in the Theorem (2), the system is contractive under the defintion III.1 and $\delta$ISS under the definition III.2 with
@@ IV CLOSED-FORM APPROXIMATION
However, solving this integral is not straightforward due to the non-linearity in $F$. We aim to find an approximate form of (15) that keeps the same dynamic properties of the original system. Given the following quantities
$\small\begin{split}f=&-D_{x}f\operatorname*{diag}(\bm{x(T)})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S(0)}^{k}+\sum_{k=1}^{K}\bm{A_{k}}^{T}\otimes\bm{S(0)}^{k}\\ f_{x}=&\rho(\hat{A}_{S}(\bm{x(T)})+\bm{b_{x}})\\ f_{\sigma}=&\rho(\hat{B}_{S}(\bm{u(0)})+\bm{b_{u}})+\rho(\hat{A}_{S}(\bm{x(0)})+\bm{b_{x}})\\ \end{split}$
@@
$\small\begin{split}\bm{x(T)|_{:}}=&e^{-(\operatorname*{diag}(\bm{b}|_{:})+\operatorname*{diag}(f_{x}|_{:})+f)T}\operatorname*{diag}(\sigma(2f_{\sigma})|_{:})\bm{x(0)|_{:}}\\ &+\operatorname*{diag}(1-\sigma(2f_{\sigma})|_{:})\sigma_{u}|_{:}.\end{split}$ (16)
Reminding that $\sigma_{u}=\sigma_{c}(B_{S}(\bm{u(0)}))$, one can easily demonstrate that $\bm{x(T)}\in\bm{\mathcal{X}}$ since its expression is a convex combination with coefficient $\sigma(2f_{\sigma})$ and $1-\sigma(2f_{\sigma})$. This closed-form solution maintains the contraction rate as stated by the following theorem.
@@
With $t=[0,T]$, the state trajectory (16) is a closed-form approximation of the dynamic system (8) with the same contraction rate $c$ under the constraints of Theorem 2.
@@
We start the proof by highlighting the relationship between the superior of the induced log-norm and the superior of the time derivative of the induced norm of the matrix exponential
@@
Therefore, we can show that the equivalence between the right-hand side of the previous relation and the induced norm of the jacobian ($D_{x}\bm{x(T)}$) of (16). Let’s name $F_{ex}$ the argument of the exponential in the (16). $||D_{x}\bm{x(t)}||_{\infty}$ satisfies the following :
$\small\begin{split}||D_{x}\bm{x(t)|_{:}}||_{\infty}&=||e^{F_{ex}}\operatorname*{diag}(\sigma(2f_{\sigma})|_{:})+\\ &\operatorname*{diag}(x|_{:})e^{F_{ex}}(\operatorname*{diag}(\sigma(2f_{\sigma})|_{:})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}t+\\ &\operatorname*{diag}(x|_{:})e^{F_{ex}}2\operatorname*{diag}(D_{\sigma}|_{:})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}-\\ &\operatorname*{diag}(\sigma_{u})2\operatorname*{diag}(D_{\sigma}|_{:})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}||_{\infty}\\ &\leq||e^{F_{ex}}||_{\infty}+||e^{F_{ex}}\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}t||_{\infty}+\\ &0.5(||e^{F_{ex}}||_{\infty}+1)||\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}||_{\infty}\\ \leq||e^{F_{ex}}||_{\infty}+&||e^{F_{ex}}\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}t||_{\infty}+||\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}||_{\infty}\end{split}$ (18)
where the last inequalities come from reasoning similar to the proof in the theorem 2 and the Lipschitz constant of $\sigma(2x)$ equal to $0.5$. The following time derivative proves the equality (17)
$\small\begin{split}\frac{\partial}{\partial t}\sup\limits_{x}||D_{x}\bm{x(t)|_{:}}||_{\infty}\big{|}_{t=0^{+}}=&-||\bm{b}||_{\infty}-||A_{1,K}||_{\infty}||\tilde{S}_{1,K}||_{\infty}-\\ &||\bm{b_{x}}||_{\infty}+||\hat{A}_{0,K}^{T}||_{\infty}||\bar{S}_{0,K}||_{\infty}\end{split}$ (19)
@@
With a fixed $T$, the state trajectory (16) can serve as a discrete-time system that in training and evaluation requires one iteration/communication to be evaluated, against the $n$ iteration required by the continuous system. However, equation (16) can not be directly distributed over the agents due to the kronacker product terms appearing in $f$. Moreover, notably, the exponential term in recursion causes vanishing gradient problems in training. Therefore, we modified the trajectory (16) to tackle these two issues. First, we multiply the kronacker product terms by $\bm{x}/(\bm{x}+\epsilon)$, with /// being the division element by element and a small $\epsilon$. Second, we replace the exponential term with a sigmoidal nonlinearity which decreases much more smoothly. The resulting closed-form graph time constant (CfGC) system is:
$\small\begin{cases}f_{\sigma}=\rho(\hat{B}_{S}(\bm{u})+\bm{b_{u}})+\rho(\hat{A}_{S}(\bm{x})+\bm{b_{x}})\\ f_{x}=\rho(\hat{A}_{S}(\bm{x})+\bm{b_{x}})\\ f=-D_{x}f\hat{A}_{S}(\bm{x})+\sum_{k=1}^{K}\bm{S}^{k}\bm{x}\bm{A_{k}}/(\bm{x}+\epsilon)\\ \bm{x}^{+}=(\bm{x}\circ\sigma(-(\bm{b}+f_{x}+f)t+\pi)-\sigma_{u})\circ\sigma(2f_{\sigma})+\sigma_{u}\end{cases}$ (20)

### v3

@@ I INTRODUCTION
Communication is a crucial element in achieving distributed solutions for multi-agent systems (MAS) from control to planning [2]. Like many distributed control algorithms, also learning-based control can benefit from inter-agent communication to partition the prediction process across multiple machines contributing to task scalability and prediction accuracy [6].
@@
In this study, we introduce a continuous graph ordinary differential equation (ODE) network called Liquid-Graph Time-Constant (LGTC) network, inspired by its single-agent counterpart, the Liquid Time Constant (LTC) network [24]. The LTC network demonstrates a strong representation learning capability, empowering agents to extrapolate and make inferences even in unfamiliar scenarios. Additionally, its closed-form version [25] eliminates the need to solve the initial value problem of the ODE, a requirement in other ODE networks [26, 27, 28]. In the existing literature, there are few attempts to make graph ODEs for modelling dynamic network systems [29], predicting traffic flow [30], or for sequential recommendations [31]. These models typically employ an auto-regressive network over a vector field modelled by a Graph Neural Network (GNN), initializing the internal state with the previous layer’s output. In contrast, the proposed LGTC aims to introduce a novel model that learns a dynamic (or “liquid”) time constant, influenced by both input data and the internal state communicated over the agent graph. This dynamic time constant enables each element of the state to capture specific dynamics, thereby enhancing the model’s predictive capabilities. The specialized dynamics can be used to reduce the number of the network’s internal state (memory) and consequently the number of information each agent needs to communicate. The final aim of this study is to approach model-based algorithms’ communication efficiency while increasing prediction accuracy, which in control translates into a higher distributed control efficiency.
@@ II PRELIMINARIES
where $\bm{s}_{i}$ is the $i$-th row of $S$. The signal manipulation can be operated by means of any graph shift operator $S\subseteq\mathcal{S}$, e.g., Laplacian, adjacency matrix, weighted Laplacian , which respects the sparsity pattern of the graph. Later, in Sect. V, we will use the Laplacian as support matrix, as it is commonly used in distributed control. However, the proposed techniques do not assume the use of a specific support matrix.
Performing $k$ repeated applications of $S$ on the same signal represents the aggregation of the $k$-hop neighbourhood information. In analogy with traditional signal processing, this property can be used to define a linear graph filtering [32] that processes the multi features signal $\bm{x}\in\mathbb{R}^{N\times G}$ with $G$ features:
$\small H_{S}(\bm{x})=\sum_{k=0}^{K}S^{k}\bm{x}\bm{H_{k}}$ (2)
where the weights $\bm{H_{k}}\in\mathbb{R}^{G\times F}$ define the output of the filter. Note that $S^{k}=S(S^{k-1})$, so that it can be computed locally with repeated 1-hop communications between a node and its neighbors. Hence, the computation of $H_{\bm{S}}$ is distributed on each node.
@@ II-B Gated Graph Neural Network
Forgetting factors can be applied to mitigate this problem, reducing the influence of past or new signal on the state. A Gated Graph Neural Network (GGNN) [23] is a recurrent Graph Neural Network that uses a gating mechanism to control how the past information influences the update of the GNN states. We can add a state and an input gates, $\bm{\hat{q}},\bm{\tilde{q}}\in Q\subseteq[0,1]^{N\times F}$, that are multiplied via the Hadamard product $\circ$ by the state and the inputs of the network, respectively. These two gates regulate how much the past information and the input are used to update the network’s internal state. GGNNs admit the following state-space representation [36],
@@
We identify the induced $\infty$-norm as $||\cdot||_{\infty}$. We used the following notation for the filters in the system (4):
@@
where $K$ is the filters length. Then, in light of assumption 1 and knowing that $||\bm{x}||_{\infty}\leq 1$, each gate feature $q_{i}$ satisfies:
$|\hat{q}_{i}|$ $\leq\sigma(||S_{I,K}||_{\infty}(||\hat{A}_{0,K}||_{\infty}+||\hat{B}_{0,K}||_{\infty})+||\hat{\bm{b}}||_{\infty})\triangleq\sigma_{\hat{q}}.$ (6)
@@ III LIQUID-GRAPH TIME CONSTANT NETWORK
In the following results on the system, we used the vector operator $\bm{X}_{|}$ that rearranges the elements of matrix $\bm{X}$ in a vector. For the system in (8), the following lemma holds
@@
$\begin{split}\dot{V}&=-\bm{x}_{|}^{T}\operatorname*{diag}(\bm{b}_{|})\bm{x}_{|}-\bm{x}_{|}^{T}\operatorname*{diag}(\bm{f}_{|})\bm{x}_{|}\\ &-\bm{x}_{|}^{T}\sum_{k=1}^{K}\bm{A_{k}}^{T}\otimes\bm{S}^{k}\bm{x}_{|}+\bm{x}_{|}^{T}\operatorname*{diag}(\bm{f}_{|})\sigma_{c}(B_{S}(\bm{u}))_{|}\end{split}$
@@
$\small\dot{V}<-\bm{x}_{|}^{T}\operatorname*{diag}(\bm{f}_{|})\bm{x}_{|}+\bm{x}_{|}^{T}\operatorname*{diag}(\bm{f}_{|})\sigma_{c}(B_{S}(\bm{u}))_{|}$
When $x_{i}>1$, $\dot{V_{i}}<0$, because the term $\bm{x}_{|}^{T}\operatorname*{diag}(\bm{f}_{|})\sigma_{c}(B_{S}(\bm{u}))_{|}\leq\bm{1}^{T}\operatorname*{diag}(\bm{f}_{|})\bm{1}$ and reminding that $f_{i}>0$. The proof is completed by noting that $\dot{V_{i}}<0$ when $x_{i}<-1$ with a similar reasoning. ∎
We denote the vector field of the dynamic system (8) with $F(x,u,S,t):\mathcal{X}\times\mathcal{U}\times\mathcal{S}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{N\times F}$ and its jacobian with $D_{x}F=\frac{\partial F_{|}}{\partial\bm{x}_{|}}$. Giving the induced infinite log-norm definition $\mu_{\infty}(\bm{X})=\max_{i}(x_{ii}+\sum_{j=1,j\neq 1}^{n}|x_{ij}|)$ [37], the followings results are known [15]:
@@
The vector field $F$ is strongly infinitesimally contracting if:
@@
For the neural network (8), we provide the following theorem
@@
$\small\begin{split}\sup\limits_{x}\mu_{\infty}(D_{x}F)=&\sup\limits_{x}\mu_{\infty}(-\operatorname*{diag}(\bm{b_{|}})-\operatorname*{diag}(\bm{f}_{|})\\ -\sum_{k=1}^{K}\bm{A_{k}}^{T}\otimes\bm{S}^{k}+\operatorname*{diag}(&\sigma_{u|}-\bm{x}_{|})\operatorname*{diag}(D_{x}\bm{f})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k})\\ \leq-||\bm{b}||_{\infty}+\mu_{\infty}(&-\sum_{k=1}^{K}\bm{A_{k}}^{T}\otimes\bm{S}^{k})+\sup\limits_{x}[-||\bm{f}_{|}||_{\infty}\\ +\mu_{\infty}(\operatorname*{diag}(&\sigma_{u|}-\bm{x}_{|})\operatorname*{diag}(D_{x}\bm{f})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k})]\end{split}$
where $D_{x}\bm{f}\in[0,1]^{NF}$ is the derivative of $\bm{f}$ in $\bm{x}_{|}$. The last inequality follows by $\mu_{\infty}(\bm{X})=||\bm{X}||_{\infty}$ when $\bm{X}>0$ and by the fact that $\bm{b}_{|}\geq 0$ and $\bm{f}_{|}\geq 0$. Moreover, we can note that $-||\bm{f}_{|}||_{\infty}\leq-||\rho(\hat{A}_{S}(\bm{x})+\bm{b_{x}})_{|}||_{\infty}$. Giving $\sigma_{u|}\in[-1,1]^{NF}$ and $-\mu_{\infty}(\bm{A_{k}}^{T}\otimes\bm{S}^{k})\leq 0$, by lemma 1, $\bm{x}_{|}\in[-1,1]^{NF}$. Therefore, we get
@@
From the statement in the Theorem 2, the system is contractive under the defintion III.1 and $\delta$ISS under the definition III.2 with
@@ IV CLOSED-FORM APPROXIMATION
However, solving this integral is not straightforward due to the non-linearity in $F$. We aim to find an approximate form of eq. (15) that keeps the same dynamic properties of the original system. Given the following quantities
$\small\begin{split}\bm{f}_{i}=&-D_{x}\bm{f}\operatorname*{diag}(\bm{x(T)}_{|})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S(0)}^{k}+\sum_{k=1}^{K}\bm{A_{k}}^{T}\otimes\bm{S(0)}^{k}\\ \bm{f}_{x}=&\rho(\hat{A}_{S}(\bm{x(T)})+\bm{b_{x}})\\ \bm{f}_{\sigma}=&\rho(\hat{B}_{S}(\bm{u(0)})+\bm{b_{u}})+\rho(\hat{A}_{S}(\bm{x(0)})+\bm{b_{x}})\\ \end{split}$
@@
$\small\begin{split}\bm{x(T)_{|}}=&e^{-(\operatorname*{diag}(\bm{b}_{|})+\operatorname*{diag}(\bm{f}_{x|})+\bm{f}_{i})T}\operatorname*{diag}(\sigma(2\bm{f}_{\sigma})_{|})\bm{x(0)_{|}}\\ &+\operatorname*{diag}(1-\sigma(2\bm{f}_{\sigma})_{|})\sigma_{u|}.\end{split}$ (16)
Reminding that $\sigma_{u}=\sigma_{c}(B_{S}(\bm{u(0)}))$, one can easily demonstrate that $\bm{x(T)}\in\bm{\mathcal{X}}$ since its expression is a convex combination with coefficient $\sigma(2\bm{f}_{\sigma})$ and $1-\sigma(2\bm{f}_{\sigma})$. This closed-form solution maintains the contraction rate as stated by the following theorem.
@@
With $t=[0,T]$, the state trajectory (16) is a closed-form approximation of the dynamic system (8) with the same contraction rate $c$ of the Theorem 2.
@@
The theorem can be demonstrated by showing that the closed-loop system maintains the same dynamic properties. We start the proof by highlighting the relationship between the superior of the induced log-norm and the superior of the time derivative of the induced norm of the matrix exponential
@@
Therefore, we can show that the equivalence between the right-hand side of the previous relation and the induced norm of the jacobian ($D_{x}\bm{x(T)}$) of eq. (16). Let’s name $F_{ex}$ the argument of the exponential in the eq. (16). $||D_{x}\bm{x(t)}||_{\infty}$ satisfies the following :
$\small\begin{split}||D_{x}\bm{x(t)_{|}}||_{\infty}&=||e^{F_{ex}}\operatorname*{diag}(\sigma(2f_{\sigma})_{|})+\\ &\operatorname*{diag}(\bm{x}_{|})e^{F_{ex}}(\operatorname*{diag}(\sigma(2f_{\sigma})_{|})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}t+\\ &\operatorname*{diag}(\bm{x}_{|})e^{F_{ex}}2\operatorname*{diag}(D_{\sigma|})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}-\\ &\operatorname*{diag}(\sigma_{u})2\operatorname*{diag}(D_{\sigma|})\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}||_{\infty}\\ &\leq||e^{F_{ex}}||_{\infty}+||e^{F_{ex}}\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}t||_{\infty}+\\ &0.5(||e^{F_{ex}}||_{\infty}+1)||\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}||_{\infty}\\ \leq||e^{F_{ex}}||_{\infty}+&||e^{F_{ex}}\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}t||_{\infty}+||\sum_{k=0}^{K}\bm{\hat{A}_{k}}^{T}\otimes\bm{S}^{k}||_{\infty}\end{split}$ (18)
where the last inequalities come from similar reasons to the proof in the theorem 2 and the Lipschitz constant of $\sigma(2x)$ equal to $0.5$. The following time derivative proves the equality (17)
$\small\begin{split}\frac{\partial}{\partial t}\sup\limits_{x}||D_{x}\bm{x(t)_{|}}||_{\infty}\big{|}_{t=0^{+}}=&-||\bm{b}||_{\infty}-||A_{1,K}||_{\infty}||\tilde{S}_{1,K}||_{\infty}-\\ &||\bm{b_{x}}||_{\infty}+||\hat{A}_{0,K}^{T}||_{\infty}||\bar{S}_{0,K}||_{\infty}\end{split}$ (19)
@@
With a fixed $T$, the state trajectory (16) can serve as a discrete-time system that in training and evaluation requires one iteration/communication to be evaluated, against the $n$ iteration required by the continuous system. However, equation (16) can not be directly distributed over the agents due to the kronacker product terms appearing in $f_{i}$. Moreover, notably, the exponential term in recursion causes vanishing gradient problems in training. Therefore, we modified the trajectory (16) to tackle these two issues. First, we multiply the kronacker product terms by $\bm{x}/(\bm{x}+\epsilon)$, with /// being the division element by element and a small $\epsilon$. Second, we replace the exponential term with a sigmoidal nonlinearity which decreases much more smoothly. The resulting closed-form graph time constant (CfGC) system is:
$\small\begin{cases}f_{\sigma}=\rho(\hat{B}_{S}(\bm{u})+\bm{b_{u}})+\rho(\hat{A}_{S}(\bm{x})+\bm{b_{x}})\\ f_{x}=\rho(\hat{A}_{S}(\bm{x})+\bm{b_{x}})\\ f_{i}=-D_{x}f\hat{A}_{S}(\bm{x})+\sum_{k=1}^{K}\bm{S}^{k}\bm{x}\bm{A_{k}}/(\bm{x}+\epsilon)\\ \bm{x}^{+}=(\bm{x}\circ\sigma(-(\bm{b}+f_{x}+f_{i})t+\pi)-\sigma_{u})\circ\sigma(2f_{\sigma})+\sigma_{u}\end{cases}$ (20)